I need help with this question, I have a few of the answers but need the rest.
Thanks Jeff
1) Use the arithmetic series of numbers 1, 3, 5, 7, 9,…to find the following:
a) What is d, the difference between any 2 terms?
Answer: 2
Show work in this space.
1+2=3+2=5+2=7+2=9
b) Using the formula for the nth term of an arithmetic series, what is 101st term? Answer:
Show work in this space.


c) Using the formula for the sum of an arithmetic series, what is the sum of the first 20 terms? 400
d) Answer:Show work in this space




e) Using the formula for the sum of an arithmetic series, what is the sum of the first 30 terms?
Answer:
Show work in this space




Answer by stanbon(75887)

The 1st term is a(1)=1
d is the difference between any two consecutive terms: e.g. d=3-1=2
The nth term is a(n)=a(1)+(n-1)d
The sum of n terms is S(n)=(1/2)n[a(1)+a(n)]
101st term: a(101)=a(1)+(100)d
=1+100*2=201
Sum of 1st 20 terms:
a(20)=1+(19)2=1+38=39
S(20)=(1/2)(20)(1+39)
=10(40)=400
Sum of 1st 30 terms:
a(30)=a(1)+29(2)
=1+58=59
S(30)=(1/2)(30)(1+59)
=15(60)
=900
Cheers,
stan H.